OpenStudy (anonymous):

by the end of the summer, sarah will earn 420 doing yard work if she deposits this amount in a savings account paying 3.5% simple annual intrest what will her bvalance be at the end of 5 years

OpenStudy (anonymous):

493.50

OpenStudy (nincompoop):

\[A=P \left( 1+\frac{ r }{ n } \right)^{nt}\] where "A" is the ending amount, "P" is the beginning amount (or "principal"), "r" is the interest rate (expressed as a decimal), "n" is the number of compoundings a year, and "t" is the total number of years. Regarding the variables, n refers to the number of compoundings in any one year, not to the total number of compoundings over the life of the investment. If it is compounded yearly, then n = 1; if it is compounded bi-annually, then n = 2; quarterly, n = 3… etc

OpenStudy (nincompoop):

so when applying this formula we have: \[A = 420\left( 1+\frac{ 0.035 }{ 1 } \right)^{1 \times 5}\]

OpenStudy (nincompoop):

but that formula is for COMPOUNDED, you are being asked for SIMPLE INTEREST meaning it is linear

OpenStudy (nincompoop):

\[I= Prt \] \[(420 \times 0.035) \times 5 = ?\] I: interest P: Principal (the money deposited or invested r: rate (in decimal form, if it is in percent divide it by 100 or move two decimal places to the left) t: time (number of years you are being asked to accrue the interest